Mr.Alavardo is 20 years older than his son.In 5 years,he will be 5 more than twice as old as his son will then be.Find their present ages.

Answer by macston(5194)

S=son's age; A=Mr.Alavardo's age=S+20
A+5=2(S+5)+5 Substitute for A and solve for S
(S+20)+5=2(S+5)+5
S+25=2S+10+5 subtract 15 from each side
S+10=2S subtract S from each side
10=S ANSWER Son's age is 10
A=S+20=10+20=30 ANSWER Mr.Alvardo is 30 years old
CHECK
30+5=2(10+5)+5
35=2(15)+5
35=30+5
35=35